OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve y/(x + 4 y)= x^6 − 9 at the point ( 1, [(−8)/33] ). m = ?

OpenStudy (anonymous):

after applying differentiation, work on dy/dx, substitute the given point, the result dy/dx is the slope m of a tangent line to the given point...

OpenStudy (anonymous):

the equation on left-hand can be solved using... \[d \left( \frac{ u }{ v } \right)=\frac{ v du - u dv }{ v^{2} }\]

OpenStudy (anonymous):

u = y and du = dy v = x+4y and dv = dx + 4 dy v^2 = (x+4y)^2 = x^2+8xy+14y^2 therefore... \[d \left( \frac{ y }{ x+4y } \right)=\frac{ (x+4y)dy-y(dx+4dy) }{ x^{2}+8xy+16y^{2} }\] \[=\frac{ xdy+4ydy-ydx-4ydy }{ x^2+8xy+16y^2 }\] \[=\frac{ xdy-ydx }{ x^2+8xy+16y^2 }\]

OpenStudy (anonymous):

for the right-hand side... \[d(x^6-9)=6x^5dx-0=6x^5dx\]

OpenStudy (anonymous):

equating both sides... \[\frac{ xdy-ydx }{ x^2+8xy+16y^2 }=6x^5dx\] \[xdy-ydx=(x^2+8xy+16y^2)6x^5dx\] \[xdy-ydx=6x^7dx+48x^6ydx+96x^5y^2dx\] \[xdy=(6x^7+48x^6y+96x^5y^2+y)dx\] \[\frac{ dy }{ dx }=\frac{ 6x^7+48x^6y+96x^5y^2+y }{ x }\]

OpenStudy (anonymous):

substitute now the coordinates of the given point... the result is the slope :-)